# Rethinking Out-of-Distribution Detection on Imbalanced Data Distribution

### v1

@@ Abstract
Detecting and rejecting unknown out-of-distribution (OOD) samples is critical for deployed neural networks to void unreliable predictions. In real-world scenarios, however, the efficacy of existing OOD detection methods is often impeded by the inherent imbalance of in-distribution (ID) data, which causes significant performance decline. Through statistical observations, we have identified two common challenges faced by different OOD detectors: misidentifying tail class ID samples as OOD, while erroneously predicting OOD samples as head class from ID. To explain this phenomenon, we introduce a generalized statistical framework, termed ImOOD, to formulate the OOD detection problem on imbalanced data distribution. Consequently, the theoretical analysis reveals that there exists a class-aware bias item between balanced and imbalanced OOD detection, which contributes to the performance gap. Building upon this finding, we present a unified training-time regularization technique to mitigate the bias and boost imbalanced OOD detectors across architecture designs. Our theoretically grounded method translates into consistent improvements on the representative CIFAR10-LT, CIFAR100-LT, and ImageNet-LT benchmarks against several state-of-the-art OOD detection approaches. Code will be made public soon.
@@ 1 Introduction
Several efforts have been applied to enhance OOD detection on imbalanced data distributions [47, 19, 41]. They mainly attribute the performance degradation to misidentifying samples from tail classes as OOD (due to the lack of training data), and concentrate on improving the discriminability for tail classes and out-of-distribution samples [47, 51]. Whereas, we argue that the confusion between tail class and OOD samples presents only one aspect of the imbalance problem arising from the long-tailed data distribution.
@@ 2 Related Work
OOD detection on imbalanced data distribution. In real-world scenarios, the deployed data frequently exhibits long-tailed distribution, and Liu et al. [30] start to study the open-set classification on class-imbalanced setup. Wang et al. [47] systematically investigate the performance degradation for OOD detection on imbalanced data distribution, and develop a partial and asymmetric contrastive learning (PASCL) technique to tackle this problem. Wei et al. [51] and [36] extend the feature-space optimization by introducing abstention classes or outlier class learning and integrating with data augmentation or margin learning techniques, respectively. Sapkota and Yu [41] employ adaptive distributively robust optimization (DRO) to quantify the sample uncertainty from imbalanced distributions. Choi et al. [8] focus on the imbalance problem in OOD data, and develop an adaptive regularization for each OOD sample during optimization. In particular, Jiang et al. [20] also utilize class prior to boost imbalanced OOD detector, which is however constrained as a heuristic post-hoc normalization for pre-trained models and unable to translate into training a better detector.
@@ 4.1 Setup
Evaluation Metrics. For OOD detection, we follow Hendrycks et al. [17], Yang et al. [52], Wang et al. [47] to report three metrics: (1) AUROC, the area under the receiver operating characteristic curve, (2) AUPR, the area under precision-recall curve., and (3) FPR95, the false positive rate of OOD samples when the true positive rate of ID samples are 95%. For ID classification, we measure the macro accuracy of the classifier. We report the mean and standard deviation of performance (%) over six random runs for each method.
@@ B.1 Training settings
For a fair comparison, we mainly follow PASCL’s [47] settings. On CIFAR10/100-LT benchmarks, we train ResNet18 [15] for 200 epochs using Adam optimizer, with a batch size of 256. The initial learning rate is 0.001, which is decayed to 0 using a cosine annealing scheduler. The weight decay is $5\times 10^{-4}$. On the ImageNet-LT benchmark, we train ResNet50 [15] for 100 epochs with SGD optimizer with the momentum of 0.9. The batch size is 256. The initial learning rate is 0.1, which is decayed by a factor of 10 at epochs 60 and 80. The weight decay is $5\times 10^{-5}$. During training, each batch contains an equal number of ID and OOD data samples (i.e., 256 ID samples and 256 OOD samples).
@@ C.2 Detailed results on CIFAR10/100-LT benchmarks
Appendix D Resource Requirement
We use 2 NVIDIA V100-32G GPUs in all our experiments.

### v2

@@ Abstract
Detecting and rejecting unknown out-of-distribution (OOD) samples is critical for deployed neural networks to void unreliable predictions. In real-world scenarios, however, the efficacy of existing OOD detection methods is often impeded by the inherent imbalance of in-distribution (ID) data, which causes significant performance decline. Through statistical observations, we have identified two common challenges faced by different OOD detectors: misidentifying tail class ID samples as OOD, while erroneously predicting OOD samples as head class from ID. To explain this phenomenon, we introduce a generalized statistical framework, termed ImOOD, to formulate the OOD detection problem on imbalanced data distribution. Consequently, the theoretical analysis reveals that there exists a class-aware bias item between balanced and imbalanced OOD detection, which contributes to the performance gap. Building upon this finding, we present a unified training-time regularization technique to mitigate the bias and boost imbalanced OOD detectors across architecture designs. Our theoretically grounded method translates into consistent improvements on the representative CIFAR10-LT, CIFAR100-LT, and ImageNet-LT benchmarks against several state-of-the-art OOD detection approaches. Code is available at https://github.com/alibaba/imood.
@@ 1 Introduction
As Wang et al. [51] reveal, a naive combination of long-tailed image cognition [39] and general OOD detection [20] techniques cannot simply mitigate this issue, and several efforts have been applied to study the joint imbalanced OOD detection problem [51, 22, 45]. They mainly attribute the performance degradation to misidentifying samples from tail classes as OOD (due to the lack of training data), and concentrate on improving the discriminability for tail classes and out-of-distribution samples [51, 55]. Whereas, we argue that the confusion between tail class and OOD samples presents only one aspect of the imbalance problem arising from the long-tailed data distribution.
@@ 2 Related Work
OOD detection on imbalanced data distribution. In real-world scenarios, the deployed data frequently exhibits long-tailed distribution, and Liu et al. [34] start to study the open-set classification on class-imbalanced setup. Wang et al. [51] systematically investigate the performance degradation for OOD detection on imbalanced data distribution, and develop a partial and asymmetric contrastive learning (PASCL) technique to tackle this problem. Consequently, Wei et al. [55] and [40] extend the feature-space optimization by introducing abstention classes or outlier class learning and integrating with data augmentation or margin learning techniques, respectively. Sapkota and Yu [45] employ adaptive distributively robust optimization (DRO) to quantify the sample uncertainty from imbalanced distributions. Choi et al. [10] focus on the imbalance problem in OOD data, and develop an adaptive regularization for each OOD sample during optimization. In particular, Jiang et al. [23] also utilize class prior to boost imbalanced OOD detector, which is however constrained as a heuristic post-hoc normalization for pre-trained models and unable to translate into training a better detector.
@@ 4.1 Setup
Evaluation Metrics. For OOD detection, we report three metrics: (1) AUROC, the area under the receiver operating characteristic curve, (2) AUPR, the area under the precision-recall curve, and (3) FPR95, the false positive rate of OOD samples when the true positive rate of ID samples are 95%. For ID classification, we measure the macro accuracy of the classifier. We report the mean and standard deviation of performance (%) over six random runs for each method.
@@ 4.2 Main Results
Additional comparison following ClassPrior’s setting. Since ClassPrior [23] uses a totally different setting against the literature [51, 40] on the ImageNet datasets, including different ID imbalance ratio and OOD test sets, we additionally compare with ClassPrior in Tab. 3. For a fair comparison, as ClassPrior does not leverage real OOD data for training, we eliminate the auxiliary ImageNet-Extra dataset and utilize the recent VOS [14] technique to generate OOD syntheses for our regularization. According to Tab. 3, our method consistently outperforms ClassPrior by a large margin on all subsets.
Table 3: Comparison on ClassPrior’s ImageNet-LT-a8 benchmark. The base model is MobileNet.
Method iNaturalist SUN Places Textures
AUROC↑ FPR95↓ AUROC↑ FPR95↓ AUROC↑ FPR95↓ AUROC↑ FPR95↓
ClassPrior 82.51 66.06 80.08 69.12 74.33 79.41 69.58 78.07
Ours 86.15 59.13 81.29 65.88 77.57 76.26 72.82 72.73
@@ 4.3 Ablation Studies
4.4 ImOOD’s Inference-time Application
Despite our main focus on training more balanced OOD detectors, we also make some attempts to apply our method during pre-trained models’ inference. According to our Theorem 3.2, for an existing OOD detector $P(i|x)$ (e.g., trained with BinDisc), we can calculate the bias term $\beta(x)$ to regulate the vanilla scorer $P(i|x)$ into balanced $P^{bal}(i|x)=\beta(x)\cdot P(i|x)$. However, as $\beta(x)=\sum_{y}{\gamma_{y}(x)\frac{P(y|x,i)}{P(y)}}$, the estimation of $\gamma_{y}(x)$ presents considerable difficulty without training, but we have also tried some trivial approaches in Tab. 7.
Table 7: Attempts to apply our ImOOD into pre-trained models’ inference stages.
Method Detector AUROC↑ AUPR↑ FPR95↓
BinDisc $P(i|x)$ 90.06 88.72 33.39
+Ours (infer) $\beta_{1}(x)P(i|x)$ 90.34 88.45 32.10
+Ours (infer) $\hat{\beta}(x)P(i|x)$ 90.86 88.95 30.80
+Ours (train) $\beta(x)P(i|x)$ 92.23 91.92 29.95
First, we simply treat $\gamma_{y}(x)$ as a constant (e.g., $\gamma_{y}(x)\equiv\gamma_{1}=1$) for arbitrary input $x$ and class $y$ to calculate the bias term (denoted as $\beta_{1}(x)$), and the results on CIFAR10-LT benchmark immediately witness a performance improvement (e.g., 0.28% increase on AUROC and 1.29% decrease on FPR95) compared to the baseline OOD detector. However, the improvement is relatively insignificant, and the phenomenon is consistent with our ablation studies in Tab. 4, which demonstrates the importance of learning a class-dependent and input-dependent $\gamma_{y}(x)$ during training.
Then, inspired by the statistical results in Fig. 2, we take a further step to use a polynomial (rank=2) to fit the curve between the predicted class $y$ and $\gamma_{y}(x)$ learned by another model, and apply the coefficients to estimate a class-dependent $\hat{\gamma}_{y}$ for the baseline model (denoted as $\hat{\beta}(x)P(i|x)$). This operation receives further enhancement on OOD detection and gets close to our learned model (e.g., 30.80% v.s. 29.95% of FPR95).
In conclusion, our attempts illustrate the potential of applying our method to an existing model without post-training, and we will continue to extend $\hat{\gamma}_{y}$ to an input-dependent version (say $\hat{\gamma}_{y}(x)$) in our future work.
@@ 5 Conclusion and Discussion
Acknowledgments and Disclosure of Funding
This work was supported in part by the Fundamental Research Funds for the Central Universities, in part by Alibaba Cloud through the Research Intern Program, and in part by Zhejiang Provincial Natural Science Foundation of China under Grant No. LDT23F01013F01.
@@ B.1 Training Settings
For a fair comparison, we mainly follow PASCL’s [51] settings. On CIFAR10/100-LT benchmarks, we train ResNet18 [18] for 200 epochs using Adam optimizer, with a batch size of 256. The initial learning rate is 0.001, which is decayed to 0 using a cosine annealing scheduler. The weight decay is $5\times 10^{-4}$. On the ImageNet-LT benchmark, we train ResNet50 [18] for 100 epochs with SGD optimizer with the momentum of 0.9. The batch size is 256. The initial learning rate is 0.1, which is decayed by a factor of 10 at epochs 60 and 80. The weight decay is $5\times 10^{-5}$. During training, each batch contains an equal number of ID and OOD data samples (i.e., 256 ID samples and 256 OOD samples). We use 2 NVIDIA V100-32G GPUs in all our experiments.
@@ C.2 Detailed Results on CIFAR10/100-LT Benchmarks
C.3 Comparison on Different Imbalance Ratios
In our manuscript, we mainly take the default imbalance ratio ($\rho=100$), which means the least frequent (tail) class only has $\frac{1}{100}$ of training samples than the most frequent (head) class). Here, we follow PASCL [51] to investigate another imbalance ratio of $\rho=50$ (relatively more balanced than $\rho=100$) on CIFAR10-LT. According to Tab. A3, our method consistently surpasses PASCL on various imbalance levels, and gains a larger enhancement (e.g., near 2.0% of AUROC) on the more imbalanced scenario with $\rho=100$, which further demonstrates our effectiveness in mitigating imbalanced OOD detection.
Table A3: Evaluation on different imbalance ratio $\rho$ for the CIFAR10-LT benchmark.
Method $\rho=100$ $\rho=50$
AUROC↑ AUPR↑ FPR95↓ AUROC↑ AUPR↑ FPR95↓
OE 89.77 87.25 34.65 93.13 91.06 24.73
PASCL 90.99 89.24 33.36 93.94 92.79 22.08
Ours 92.93 92.51 28.73 94.37 94.24 19.72
C.4 Investigation on SOTA General OOD Detection Methods
Besides the method specifically designed for imbalanced OOD detection, we also compare with several recently published detectors that aims at general (or, balanced) OOD detection:
  • NECO [2] studies the neural collapse phenomenon develops a novel post-hoc method to leverage geometric properties and principal component spaces to identify OOD data.
  • fDBD [31] investigates model’s decision boundaries and proposes to detect OOD using the feature distance to decision boundaries.
  • IDLabel [15] theoretically delineates the impact of ID labels on OOD detection, and utilizes ID labels to enhance OOD detection via representation characterizations through spectral decomposition on the graph.
  • ID-like [3] leverages the powerful vision-language model CLIP to identify challenging OOD samples/categories from the vicinity space of the ID classes to further facilitate OOD detection.
The experimental results are displayed in Tab. A4. On the CIFAR10-LT benchmark, our method surpasses the recent OOD detectors by a large margin. For imbalanced data distribution, IDLabel [15] struggles to capture the distributional difference across ID classes, while the pure post-hoc methods NECO [2] and fDBD [31] even get worse results because they fail to generalize to the scenario with highly skewed feature spaces and decision boundaries. On the ImageNet-LT benchmark, even the incorporation of the powerful CLIP model cannot well address the imbalance problem for ID-like [3], while our method can specifically and effectively facilitate imbalanced OOD detection with the help of our theoretical groundness.
Table A4: Comparison with SOTA general OOD detectors.
Benchmark Method Pub&Year AUROC↑ AUPR↑ FPR95↓
CIFAR10-LT NECO ICLR’24 85.15 82.39 40.44
fDBD ICML’24 87.90 83.07 41.98
IDLabel ICML’24 90.06 88.29 34.66
Ours - 93.55 92.83 28.52
ImageNet-LT ID-like CVPR’24 72.05 71.37 78.36
Ours - 75.84 73.19 74.96
C.5 Additional Analysis on Correctly-detected ID and OOD Samples
In Fig. 1(a), we reveal that the class labels for wrongly-identified ID samples and the class predictions for wrongly-detected OOD samples are both sensitive the to the ID class distribution prior. Here, Fig. A3 indicates the correctly-detected ID and OOD samples are insensitive to the class distribution.
Figure A3: Statistics on correctly-detected ID and OOD samples.
In particular, the per-class prediction quantity of correctly-detected OOD samples in Fig. A3 may not precisely describe the statistical distribution, as the maximum ID-class prediction probability is relatively low (e.g., $max_{y}{P(y|x,i)}=0.12$ for 10-category classification) and introduce extra noise. Therefore, we supplement the statistics of per-class prediction probability in Tab. A5, and all of the distributions for OE, Energy, and PASCL are nearly even.
Table A5: Average per-class prediction probability for correctly-detected OOD samples.
Method $cls_{1}$ $cls_{2}$ $cls_{3}$ $cls_{4}$ $cls_{5}$ $cls_{6}$ $cls_{7}$ $cls_{8}$ $cls_{9}$ $cls_{10}$
OE 0.12 0.11 0.14 0.14 0.13 0.11 0.16 0.11 0.11 0.12
Energy 0.28 0.25 0.31 0.26 0.29 0.29 0.39 0.32 0.36 0.33
PASCL 0.15 0.12 0.12 0.14 0.11 0.12 0.13 0.12 0.11 0.11